One day a hot dog cart sold 71 sausages. The polish sausages cost $5 each and the kosher sausages cost $3 each. In all $297 worth of sausages were sold. How many of each kind were sold?
Answer by mananth(16949)

Polish x Numbers
Kosher ------------ y Numbers
x+y=71 .............1
5x+3y=297 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 1
-3x-3y= -213
5x+3y=297
Add the two equations
2x = 84
/2
x= 42
plug value of x in (1)
x+y= 71
42+y=71
y=71-42
y=29